### v2

# CAIMAN: Causal Action Influence Detection for Sample-efficient Loco-manipulation

Abstract
Enabling legged robots to perform non-prehensile loco-manipulation is crucial for enhancing their versatility. Learning behaviors such as whole-body object pushing often requires sophisticated planning strategies or extensive task-specific reward shaping, especially in unstructured environments. In this work, we present CAIMAN, a practical reinforcement learning framework that encourages the agent to gain control over other entities in the environment. CAIMAN leverages causal action influence as an intrinsic motivation objective, allowing legged robots to efficiently acquire object pushing skills even under sparse task rewards. We employ a hierarchical control strategy, combining a low-level locomotion module with a high-level policy that generates task-relevant velocity commands and is trained to maximize the intrinsic reward. To estimate causal action influence, we learn the dynamics of the environment by integrating a kinematic prior with data collected during training. We empirically demonstrate CAIMAN’s superior sample efficiency and adaptability to diverse scenarios in simulation, as well as its successful transfer to real-world systems without further fine-tuning.
Keywords: Reinforcement Learning, Loco-manipulation, Intrinsic Motivation
1 Introduction
Modern legged robots continue to impress with their diverse capabilities, from traversing challenging terrains [1, 2, 3] to executing agile maneuvers such as backflips and parkour [4, 5, 6]. However, enabling these systems to interact with their environments remains an open research challenge [7, 8, 9]. A common approach is to equip legged robots with external manipulators for prehensile manipulation [10, 11, 12, 13]. Although effective, these methods are often constrained by the size and weight of objects they can successfully manipulate. In contrast, non-prehensile manipulation demands whole-body coordination and remains a non-trivial task.
Traditional approaches for whole-body manipulation explicitly model robot-object interactions, often relying on complex planning and optimization for coordinating motion [14]. These methods require accurate models of both the robot and the environment, making them less scalable to high-dimensional systems with complex, stochastic dynamics, and often limiting contact to predefined regions [15, 16]. Learning-based methods offer a more scalable alternative, improving computational efficiency and reducing dependence on precise object estimation [17, 18, 19]. However, these methods frequently rely on tedious reward shaping or learning curriculums to guide exploration and foster meaningful behavior. Furthermore, the large exploration space of loco-manipulation tasks often necessitates special treatment, such as learning from behavioral priors  [20, 21, 22] or task-agnostic explorative rewards [23, 24], to encourage meaningful engagement with the environment. Prior works have also leveraged a hierarchical structure to decouple locomotion and high-level planning [25]. Training the high-level policy, however, still requires elaborate reward design and shaping.
To alleviate these challenges, we introduce CAIMAN, a framework for training non-prehensile object pushing skills in legged robots. CAIMAN employs a hierarchical control structure that decouples high-level planning from low-level locomotion. We train robust locomotion using an existing pipeline, and we learn the high-level policy with a simple yet effective reward structure consisting of only three terms: a sparse task reward, an action regularizer, and an intrinsically motivated explorative reward. The sparse task reward provides a general signal for various pushing tasks, including ones requiring navigation through obstacles. The high-level policy is incentivized to explore and gain control over the environment via an exploratory reward based on Causal Action Influence (CAI) [26], which quantifies how much influence the agent has on the states of other entities and is computed based on environment dynamics. Instead of learning the dynamics from scratch, CAIMAN utilizes a simple pre-defined kinematic prior, and combines it with learned residual dynamics that compensate for physical interactions beyond the prior model. Using only the sparse task reward, we show that CAIMAN achieves strong learning performance and superior sample efficiency against other baselines, especially in complex scenarios with obstacles. We also demonstrate that the learned residual accurately captures complex robot-object interactions that are not modeled by the kinematic prior. Finally, we successfully transfer the learned policy to a real quadruped robot, enabling it to perform whole-body object pushing in obstacle environments.
To summarize, our contribution is three-fold: 1) a general hierarchical framework for learning whole-body object pushing with legged robots in various scenarios, including navigating through obstacles; 2) an intrinsically motivated reward based on causal influence calculated from a combination of a kinematics prior and learned residual dynamics; 3) Successful hardware validation on a real-world quadruped robot.
2 Related Work
2.1 Loco-manipulation on Legged Systems
The integration of locomotion and manipulation has gained significant attention as a promising and application-oriented research field for legged robots. Traditional model-based methods that rely on accurate representations of both robot and object to optimize trajectories [11, 14, 16, 27] have shown success in tasks such as box carrying [10, 12], but often require accurate state estimation [28, 29, 30] and struggle with scalability due to the complexity of contact modes [31, 32]. Reinforcement learning (RL) is a promising alternative that avoids explicit modeling and has been successfully applied to tasks such as soccer dribbling [33, 34], button pushing [18, 35], and door opening [19, 23, 36]. However, effective exploration may be difficult to achieve due to the large search space of robot-object interactions [17, 37]. To address exploration challenges, researchers have proposed using behavior prior [38, 39] or task-agnostic explorative rewards [23, 24] as mechanisms to guide learning. Furthermore, hierarchical frameworks [40, 41] are utilized to decompose tasks into high-level planning and low-level control, enabling effective behavior across various scenarios [42, 43].
In accordance with previous works, we also implement a hierarchical control framework for whole-body object pushing, decoupling high-level velocity planning from low-level locomotion, similarly to Jeon et al. [25]. We aim to utilize a simple reward structure, as opposed to one that requires sophisticated design effort as seen in [25], to achieve sample-efficient high-level policy training, even under complex scenarios such as object manipulation through obstacles.
2.2 Intrinsically Motivated Reinforcement Learning
Intrinsic motivation (IM) [44] plays a vital role in reinforcement learning, especially when extrinsic rewards are sparse or difficult to design. Core IM mechanisms include curiosity [45], learning progress [46], and empowerment [47], each promoting exploration and skill acquisition. Curiosity, often measured via prediction errors in learned world models [48, 49, 50], rewards agents for encountering novel states, and has been previously applied for learning loco-manipulation skills [23, 24]. Learning progress encourages agents to focus on regions in the state space with rapid improvement, supporting curriculum learning and adaptive exploration [51, 52].
Our work builds on empowerment, an information-theoretic quantity defined as the channel capacity between the agent’s actions and its sensory observations [47, 53, 54]. Recent work has connected IM to causal reasoning, aiming to improve sample efficiency and interpretability [55, 56, 57, 58, 59]. To encourage effective exploration, we employ causal action influence (CAI) [26], a measure of an agent’s ability to influence its environment and a conceptual lower bound on empowerment.
3 Preliminaries
(a) No influence of $A$ on $S^{\prime}_{2}$
(b) Influence of $A$ on $S^{\prime}_{1}$ and $S^{\prime}_{2}$
Figure 1: Illustration of the LCM (left) for two different environment situations $S=s$ (right) in the loco-manipulation task. The LCM captures the transition from $S,A$ to $S^{\prime}$, factorized into state components. While the global SCM is fully connected (dashed and continuous lines), the LCM $\mathcal{G}_{S=s}$ (continuous lines) is causally minimal. We are interested in detecting the presence of continuous orange arrows in the LCM, i.e. the influence of the action $A$ on next states $S^{\prime}$.
We model decision-making in a dynamic environment as a Markov Decision Process (MDP) [60], defined by the tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $P$ the transition kernel, $R$ the reward function, and $\gamma$ the discount factor. Following the principle of independent causal mechanisms [61], we assume that the world consists of interacting but independent entities. This induces a state space factorization $\mathcal{S}=\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{N}$ for $N$ entities, where each factor $\mathcal{S}_{i}$ represents the state of entity $i$. An MDP coupled with a policy $\pi:\mathcal{S}\mapsto\mathcal{A}$ induces a Structural Causal Model (SCM) [62] describing the resulting trajectory distribution.
Definition 1 (Structural Causal Model [62]).
An SCM is a tuple $(\mathcal{U},\mathcal{V},F,P^{u})$, where $\mathcal{U}$ are set of exogenous variables (e.g., latent randomness) sampled from $P^{u}$, $\mathcal{V}$ are set of observed variables (e.g., states, actions, rewards), $F$ is the set of structural functions capturing the causal relations, such that functions $f_{V}:\mathrm{Pa}(V)\times\mathcal{U}\rightarrow V$, with $\mathrm{Pa}(V)\subset\mathcal{V}$ denoting the set of parents of $V$, determine the value of endogenous variables $V$ for each $V\in\mathcal{V}$.
SCMs are typically visualized as directed acyclic graphs, where nodes represent variables and edges indicate causal relations, as shown in Figure 1. In our case, the SCM captures one-step transitions with variables $\mathcal{V}={S_{1},\ldots,S_{N},A,S_{1}^{\prime},\ldots,S_{N}^{\prime}}$. Due to the Markov property and flow of time, causal dependencies exist only from $(S,A)$ to $S^{\prime}$. While most entity pairs may theoretically interact (i.e., $S_{i}/A\rightarrow S^{\prime}_{j}$ for most $i,j$), interactions often become sparse when we observe a specific state configuration. We use the Local Causal Model (LCM) to capture such context-specific structure [63].
Definition 2 (Local Causal Model [63]).
Given an SCM $(\mathcal{U},\mathcal{V},F,P^{u})$ and an observation $V=v,V\subset\mathcal{V}$, the local SCM is the SCM with $F_{V=v}$ and graph $\mathcal{G}_{do(V=v)}$ obtained by pruning edges from $\mathcal{G}_{do(V=v)}$ until it is causally minimal.
In this work, we build on the insight that encouraging agents to gain control over their environment facilitates learning in loco-manipulation tasks. We do so by driving the agent toward states where it can influence other entities. We make use of a principled explicit measure of local influence, the Causal Action Influence (CAI) [26, 64] measure. CAI is a state-dependent measure of control that assesses whether an agent, through its actions, can affect other entities. Graphically, it corresponds to predicting the existence of an edge $A\rightarrow S^{\prime}_{j}$ in the LCM $\mathcal{G}_{do(S=s)}$. This influence is measured with point-wise conditional mutual information $I(S_{j}^{\prime};A\mid S=s)$.
4 Method
Figure 2: CAIMAN framework: The high-level (HL) policy generates desired base velocity commands, which are translated into joint commands by a low-level (LL) policy. We utilize a simple kinematic prior and learned residual dynamics to model the robot-object interaction in the environment while providing a CAI-based explorative bonus along with the sparse task reward.
The hierarchical framework adopted by CAIMAN is presented in Figure 2. It consists of a high-level (HL) policy that generates desired base velocity commands and a low-level policy that translates velocity commands into joint-level actions. In this work, we introduce a novel, sample-efficient approach for learning the high-level policy, while building upon an existing pipeline [65] to train robust low-level policies.
4.1 Low-level Locomotion Policy
The low-level locomotion policy $\bm{\pi}_{l}$ is trained to generate target joint positions $\bm{q}_{j,des}\in\mathbb{R}^{12}$ that track a desired base velocity command $\bm{v}_{r,des}\in\mathbb{R}^{3}$. This command, $\bm{v}_{r,des}=(v^{x}_{r,des},v^{y}_{r,des},\omega^{z}_{r,des})$, specifies linear velocities in the longitudinal ($x$) and lateral ($y$) directions and a yaw rate, all in the robot frame. The desired joint positions $\bm{q}_{j,des}$ are tracked using joint-level proportional-derivative (PD) controllers to produce the corresponding joint torques $\bm{\tau}_{j,des}\in\mathbb{R}^{12}$. The policy is trained with velocity commands sampled uniformly from a predefined range. To ensure robust sim-to-real transfer and hardware performance during contact-rich tasks like pushing, we apply domain randomization and external perturbations following [65]. Further details of the low-level observation $\bm{o}_{l}$ can be found in Appendix A.
4.2 High-level Planning Policy
The high-level policy $\bm{\pi}_{h}$ is trained to generate desired robot velocity commands $\bm{a}_{h}:=\bm{v}_{r,des}$ that achieve successful task completion in object pushing. Its observation $\bm{o}_{h}$ includes the robot’s linear and angular velocities, the poses of the robot and object, the target object position $\bm{p}_{t}=(x_{p},y_{p})\in\mathbb{R}^{2}$, and the previous action—all expressed in the world frame. In scenarios with obstacles (e.g., walls), $\bm{o}_{h}$ also includes their poses in the world frame. For a detailed description of the high-level observation, refer to Appendix A.
We use Proximal Policy Optimization (PPO) [66] as the base RL algorithm, with a simple reward function composed of three terms:
$r=w_{1}\mathds{1}_{\left\lVert\bm{p}_{o}-\bm{p}_{t}\right\rVert_{2}<\epsilon}+w_{2}r_{\text{CAI}}+w_{3}\left\lVert\bm{a}_{h}-\bm{a}_{h,prev}\right\rVert_{2}^{2},$ (1)
where $\bm{p}_{o}$, $\bm{p}_{t}$ are the current and target object positions, $\epsilon$ is a success threshold, and $\bm{a}_{h}$, $\bm{a}_{h,prev}$ are the current and previous high-level actions, respectively. The exploration bonus $r_{\text{CAI}}$ is derived from the CAI measure $\tilde{C}_{j}$ over the object $j$. We define the CAI reward as:
$r_{\text{CAI}}=\tilde{C}_{j=\text{object}}(s)=\frac{1}{K}\sum_{i=1}^{K}D_{\mathrm{KL}}\left(P_{S_{j}^{\prime}\mid S=s,A=a^{(i)}}\bigg{\|}\frac{1}{K}\sum_{k=1}^{K}P_{S_{j}^{\prime}\mid S=s,A=a^{(k)}}\right),$ (2)
given $K$ actions $\{a^{(i)}\}_{i=1}^{K}$ sampled from the policy. This approximation estimates the marginal $P_{S_{j}^{\prime}\mid s}$ using Monte Carlo sampling. We model the transition distribution $P_{S_{j}^{\prime}\mid S=s,A=a}$ as a fixed-variance Gaussian (details in Section 4.3), which enables closed-form KL divergence calculation using the Gaussian mixture approximation [67].
The CAI reward encourages the agent to reach states where it exerts greater influence over the object, thereby promoting task-relevant exploration and learning. The weight $w_{2}$ for the CAI reward scales with the raw CAI score:
$w_{2}=w_{2,b}+\max\left(0,(r_{\text{CAI}}-\alpha_{1})/\alpha_{2}\right),$ (3)
where $\alpha_{1}$ is the threshold for scaling and $\alpha_{2}$ controls the scaling rate.
Finally, we foster more directed exploration by injecting time-correlated noise into the action sampling process during training [68, 69]. For more details on training the high-level policy, we refer the reader to Appendix D.
4.3 Dynamics learning
To calculate the exploration bonus $r_{\text{CAI}}$ in (2), we learn the transition model $P_{S_{j}^{\prime}\mid S=s,A=a^{(k)}}$ for all entities $j$. In our current setting, we focus on a single entity—the pushable object—but the framework naturally extends to multiple entities. Concretely, this reduces to learning the object transition model $P_{\bm{s}^{\prime}_{o}\mid\bm{s}_{h},\bm{a}_{h}}$. The object state $\bm{s}_{o}$ is defined as the object’s position, while the high-level state $\bm{s}_{h}=(\bm{\xi}_{r},\bm{\xi}_{o},\bm{v}_{r},\bm{v}_{o})$ includes the robot’s pose $\bm{\xi}_{r}$ and velocity $\bm{v}_{r}$ and the object’s pose $\bm{\xi}_{o}$ and velocity $\bm{v}_{o}$. The high-level action $\bm{a}_{h}$ corresponds to the desired robot velocity, and the next object state $\bm{s}^{\prime}_{o}=\bm{p}^{\prime}_{o}=(x^{\prime}_{o},y^{\prime}_{o})$ denotes its position at the subsequent timestep. Instead of using a pretrained model, CAIMAN leverages data collected from high-level interactions during training to efficiently learn the dynamics. As described earlier, we model the object’s transition probability $P_{\bm{s}^{\prime}_{o}\mid\bm{s}_{h},\bm{a}_{h}}$ as a fixed-variance Gaussian distribution $\mathcal{N}(\bm{s^{\prime}}_{o};f_{\theta}(\bm{s}_{h},\bm{a}_{h}),\sigma^{2})$, where $f_{\theta}$ is a neural network that predicts the mean of the distribution.
To enhance learning efficiency, we incorporate a simple kinematic prior model $f_{\text{kin}}$, which estimates the next object position $\tilde{\bm{p}}^{\prime}_{o}$ using geometric reasoning based on the relative pose between robot and object and the commanded velocity. This estimate is computed by projecting the robot’s velocity $\bm{a}_{h}$ onto the direction from the robot to the object and updating the object’s position accordingly:
$\tilde{\bm{p}}^{\prime}_{o}=\begin{cases}\bm{p}_{o}+\delta t\cdot(\bm{a}_{h,xy}\cdot\delta\hat{\bm{p}})\cdot\delta\hat{\bm{p}},\quad&\text{if }\left\lVert\bm{p}_{o}-\bm{p}_{r}\right\rVert_{2}\leq\epsilon_{p}\text{ and }\bm{a}_{h,xy}\cdot\delta\hat{\bm{p}}>0\\ \bm{p}_{o},&\text{otherwise.}\end{cases}$ (4)
Here, $\delta\hat{\bm{p}}=(\bm{p}_{o}-\bm{p}_{r})/\left\lVert\bm{p}_{o}-\bm{p}_{r}\right\rVert_{2}$ is the unit vector pointing from the robot to the object, $\delta t$ is the high-level control step size, and $\epsilon_{p}$ is a distance threshold. The conditions ensure that the robot is close enough to and moving toward the object.
We combine the kinematic prior with a learned residual model $f_{\text{res}}$, parameterized by $\theta$, to capture complex physical interactions beyond the kinematics model, including nonlinear effects such as friction, drag, and collisions with obstacles. The final dynamics model is thus:
$f_{\theta}(\bm{s}_{h},\bm{a}_{h})=f_{\text{kin}}(\bm{s}_{h},\bm{a}_{h})+f_{\text{res}}(\bm{s}_{h},\bm{a}_{h};\theta),$ (5)
where $f_{\text{kin}}$ is deterministic and independent of $\theta$. We train the residual model by minimizing the mean squared error between the predicted and true object positions:
$\mathcal{L}_{\text{res}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\lVert\tilde{\bm{s}}^{\prime(i)}_{o}-\bm{s}^{\prime(i)}_{o}\right\rVert_{2}^{2},$ (6)
where $\tilde{\bm{s}}^{\prime(i)}_{o}=f_{\text{kin}}(\bm{s}^{(i)}_{h},\bm{a}^{(i)}_{h})+f_{\text{res}}(\bm{s}^{(i)}_{h},\bm{a}^{(i)}_{h};\theta)$.
Although the high-level action $\bm{a}_{h}$ could, in principle, be sampled from the full support of the desired velocity $\bm{v}_{r,des}$ [26], not all commands are feasible for the low-level controller to execute under the robot’s current state. Therefore, we define $\bm{a}_{h}=\tilde{\bm{v}}_{r}$ to be the achievable velocity, and use $\tilde{\bm{v}}_{r}$ for both dynamics learning and CAI computation. Given training samples $\mathcal{D}={(\bm{s}_{h}^{(i)},\bm{a}_{h}^{(i)},\bm{s}^{\prime(i)}_{h})}$, we have access to the achieved robot velocity $\bm{v}^{\prime}_{r}$ at the next timestep. When computing CAI as an exploration reward, we assume that the robot’s velocity can only change within a limited range over one high-level step. Thus, for any state $\bm{s}_{h}$, we sample the action $\bm{a}_{h}$ from a bounded range centered on the current velocity:
$\bm{a}_{h}=\tilde{\bm{v}}_{r}\sim\mathcal{U}[\bm{v}_{r}\pm\delta\bm{v}_{r}]$ (7)
where $\delta\bm{v}_{r}$ is a fixed velocity deviation range. Detailed values for the velocity limits $(\delta v_{r}^{x},\delta v_{r}^{y},\delta\omega_{r}^{z})$ are provided in Appendix D.
5 Experiments
Figure 3: Illustrations for the Single-object (left), Single-wall (middle), and Multi-wall (right) tasks. The yellow sphere denotes the object’s target position.
We train both the high-level and low-level policies using Isaac Sim [70]. Training hyperparameters and additional details are provided in Appendix D. We evaluate CAIMAN on three pushing tasks of increasing difficulty, illustrated in Figure 3. For the Single-object task, the robot must manipulate a single cuboid object to a fixed target position. The Single-wall task includes a fixed wall that blocks the direct path between the object and the target. The robot must navigate the object around the wall to reach the goal. Finally, the Multi-wall task includes multiple fixed walls that obstruct the path between the object and the target. The robot must maneuver the object through these obstacles to complete the task. The presence of fixed wall obstacles in the single-wall and multi-wall tasks introduces regions in the state space where object dynamics become more complex, reducing the robot’s ability to exert consistent influence over the object. For all tasks, we train with a fixed target position. A full description of scene configurations is provided in Appendix B.
Baselines
We compare our approach against several baselines. The Heuristics baseline trains the high-level policy using a distance-based heuristic reward, $r_{\text{heu}}=\exp\left(-\left\lVert\bm{p}_{r}-\bm{p}_{o}\right\rVert_{2}\right)$, encouraging the robot to minimize its distance to the object. This mirrors the main explorative reward from prior work [25]. We also conduct comparisons with other intrinsic motivation algorithms, namely Random Network Distillation (RND) [71]. We implement two variants: RND-full, which mirrors the original RND implementation and computes the prediction error for the full state, and RND-object, which computes the prediction error for the object’s state only. As for ablations, we implement CAI-kinematic, where CAI is computed using only the kinematic prior (Equation 4), and and CAI-learned, where CAI is computed using a fully learned dynamics model (no prior).
For detailed formulations of all reward functions and their coefficients, see Appendix C.
5.1 Simulation results
HeuristicsCAI-kinematicCAI-learnedRND-fullRND-objectCAIMAN
Figure 4: Policy success rate evaluated at every 100 training iterations for all methods and tasks. Results are evaluated across 800 episodes and averaged over 3 seeds, shaded area represents standard deviation.
We evaluate the learning performance of each method by measuring the success rate for each task. A task is considered successfully completed if the distance between the object and the target is below a threshold $\epsilon_{s}$ at the end of an episode. The results are shown in Figure 4.
With only sparse task rewards, task-relevant skill acquisition heavily relies on effective exploration. In the Single-object task, all methods—except RND-full—achieve some degree of success. CAIMAN, RND-object, and Heuristics perform best, while CAI-kinematic and CAI-learned underperform. The low performance of CAI-learned is likely due to the high data and time demands required to learn an accurate dynamics model from scratch. In contrast, CAI-kinematic performs comparably to the top methods, suggesting that the naive kinematic prior is sufficient for simpler tasks with straightforward object interactions.
In the more complex Single-wall and Multi-wall tasks, only CAIMAN and RND-object demonstrate notable success, with RND-object exhibiting lower asymptotic performance than CAIMAN in both scenarios. This highlights CAIMAN’s unique ability to guide task-relevant exploration even in the presence of obstacles, validating its effectiveness in cluttered environments without relying on dense extrinsic rewards. The failure of CAI-kinematic and CAI-learned in these tasks emphasizes the importance of combining a kinematic prior with a learned residual model to efficiently capture the complex object dynamics needed for effective exploration.
RND-full results in agents randomly exploring the state space without yielding meaningful behavior in all tasks. This finding reinforces the need for object-centric exploration bonuses when the skill to be learned involves physical interaction with objects. The performance gap between RND-object and CAIMAN can be partially attributed to the phenomenon of detachment [72], where the agent drifts away from reward-depleted areas and continuously pursues new regions with higher intrinsic reward, potentially losing avenues of task-relevant exploration.
5.2 Leveraging pretrained dynamics for new tasks
Figure 5: Single wall task with target positions randomly sampled within an area in front of the wall.
Figure 6: Learning the single wall random target task using learning from scratch and learning with models transferred from the fixed target task.
Since our framework learns the environment dynamics, we propose that the dynamics residual obtained from a previous training can be reused across different tasks—as long as the underlying dynamics remain consistent. We hypothesize that leveraging a pretrained model improves the accuracy of CAI estimates in the early stages of training, thereby enhancing the exploration signal and significantly boosting sample efficiency.
To test this hypothesis, we consider a generalization of the Single-wall task in which the target position is randomized instead of fixed, sampled uniformly within a predefined area as illustrated in Figure 6. Detailed scene configurations are provided in Appendix B. We compare the original CAIMAN, which learns the dynamics residual from scratch, to one that reuses a pretrained residual model. For the latter, we initialize the residual with one from the original Single-wall task and continuously fine-tune it during training with data collected from the new generalized task.
As shown in Figure 6, the reuse of learned dynamics significantly accelerates learning with sparse rewards. The pretrained model yields more informative CAI estimates, guiding the robot toward meaningful interaction behaviors early in training. This leads to substantial gains in sample efficiency, supporting our hypothesis and demonstrating the benefits of transferring learned dynamics across related tasks.
5.3 Hardware deployment
Figure 7: Snapshots from the hardware deployment for the single object and single wall tasks. The mass of the box is 5.5 kilograms with a dimension of $(0.55,0.55,0.5)$ meters.
We validated our trained policies on real-world quadruped pushing tasks with Unitree Go2, as shown in Figure 7. To bridge the sim-to-real gap, we applied domain randomization to object mass and friction during high-level policy training. Additional details are provided in Appendix E. All entities in the environment were tracked using an external motion capture system. The trained policies are deployed directly to the robot and successfully execute pushing tasks without requiring additional fine-tuning. We refer interested readers to the supplementary video for more details.
6 Conclusion
We present CAIMAN, a general framework for training whole-body pushing skills in legged robots. CAIMAN adopts a hierarchical control strategy that decouples locomotion from high-level planning. By introducing an intrinsic exploration bonus based on CAI, our method encourages the robot to gain control over relevant entities in its environment under sparse task signals—eliminating the need for tedious, hand-crafted reward shaping, even in cluttered scenarios with obstacles. To further enhance training efficiency, we bootstrap the CAI computation using a simple yet effective kinematic prior, and refine it with a learned residual dynamics model. We evaluate CAIMAN on a suite of quadruped pushing tasks and show that it consistently outperforms competitive baselines in terms of sample efficiency. We also demonstrate that the learned dynamics residual can be reused to accelerate training for new tasks, highlighting the transferability of our approach. Finally, we validate our method on real hardware, seamlessly deploying the trained policies without the need for fine-tuning. Overall, CAIMAN offers a robust and scalable solution for learning physically grounded manipulation behaviors in legged robots, paving the way toward more autonomous and adaptable robotic systems.
7 Limitations
Although CAI-based exploration provides meaningful guidance for robot-object interaction, it inevitably relies heavily on an accurate dynamics model. Currently, CAIMAN trains the high-level policy alongside the dynamics residual, which may lead to undesired exploratory behaviors in the early stages of training when the dynamics model is still inaccurate. This can be mitigated by leveraging a pretrained world model. Our deployment pipeline currently relies on an external motion capture system. We leave the incorporation of an onboard perceptive module for future work. Another potential direction for the future is to extend CAIMAN to enable exploration for multi-object interactions.
Acknowledgments
We thank Dongho Kang, Taerim Yoon, Zijun Hui, Flavio De Vincenti and Hélène Stefanelli for developing the hardware testing framework.
Appendix A Detailed Observation Space
Below we list the observation space for low-level, high-level policies and the high-level state for calculating CAI.
Low-level observation $\bm{o}_{l}$
${}_{\mathcal{B}}\bm{v}_{r}\in\mathbb{R}^{3}$ robot linear velocity in base frame $\mathcal{B}$
${}_{\mathcal{B}}\bm{\omega}_{r}\in\mathbb{R}^{3}$ robot angular velocity in base frame $\mathcal{B}$
$\bm{q}_{j}\in\mathbb{R}^{12}$ Joint positions
$\dot{\bm{q}}_{j}\in\mathbb{R}^{12}$ Joint velocities
${}_{\mathcal{B}}\bm{g}\in\mathbb{R}^{3}$ Projected gravity in base frame $\mathcal{B}$
$\bm{v}_{r,des}=(v^{x}_{r,des},v^{y}_{r,des},\omega^{z}_{r,des})\in\mathbb{R}^{3}$ Desired velocity command
$\bm{a}_{l,prev}\in\mathbb{R}^{12}$ Previous action
High-level observation $\bm{o}_{h}$ (in world frame $\mathcal{W})$
$\bm{v}_{r}\in\mathbb{R}^{3}$ Robot linear velocity
$\bm{\omega}_{r}\in\mathbb{R}^{3}$ Robot angular velocity
$\bm{\xi}_{r}=(x_{r},y_{r},\psi_{r})\in\mathbb{R}^{3}$ Robot pose
$\bm{\xi}_{o}=(x_{o},y_{o},\psi_{o})\in\mathbb{R}^{3}$ Object pose
$\bm{p}_{t}=(x_{t},y_{t})\in\mathbb{R}^{2}$ Target position
$\bm{a}_{h,prev}\in\mathbb{R}^{3}$ Previous action
additional:
$(x_{w},y_{w},\psi_{w})\in\mathbb{R}^{3}$ Wall pose (for each wall)
high-level state for CAI $\bm{s}_{h}$ (in world frame $\mathcal{W}$)
$\bm{\xi}_{r}=(x_{r},y_{r},\psi_{r})\in\mathbb{R}^{3}$ Robot pose
$\bm{\xi}_{o}=(x_{o},y_{o},\psi_{o})\in\mathbb{R}^{3}$ Object pose
$\bm{v}_{r}=(v_{r}^{x},v_{r}^{y},\omega_{r}^{z})\in\mathbb{R}^{3}$ Robot velocity
$\bm{v}_{o}=(v_{o}^{x},v_{o}^{y},v_{o}^{z})\in\mathbb{R}^{3}$ Object velocity
Table A.1: Detailed observation space for each module.
Appendix B Scene Configurations
The detailed initial pose for each entity and target position are shown in Table B.1. In simulation, the object measures $0.5$ meters in length, width, and height, and weighs $5.0$ kilograms. The wall in the single-wall task has dimensions $(0.1,1.0,0.5)$ meters, while walls in the multi-wall task measure $(0.1,1.25,0.5)$ meters. For the transfer task of Single-wall, we sample the target position within the range of ($\mathcal{U}[3.0,4.0]$, $\mathcal{U}[-1.0,1.0]$).
Task Single-object Single-wall Multi-wall
Robot (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$) (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$) (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$)
Object (1.0, 0.0, 0.0) (1.0, 0.0, 0.0) (1.0, 0.0, 0.0)
Target (3.0, 0.0) (3.5, 0.0) (4.0, 0.0)
Wall 1 - (2.0, 0.0, 0.0) (2.0, 0.0, 0.0)
Wall 2 - - (3.25, 1.25, 0.0)
Wall 3 - - (3.25, -1.25, 0.0)
Table B.1: We list the scene configurations, where the initial pose $(x,y,\psi)$ for every existing entity and the target position $(x,y)$ for each task (Single-object, Single-wall, Multi-wall) is shown.
Appendix C Rewards and Coefficients
We list the reward functions and corresponding reward parameters for all tasks and training methods in Table C.1. The RND reward term is defined as $r_{RND}=\lVert f(s)-\hat{f}(s)\rVert_{2}$, where $s$ is the state to be encoded, and $f(s)$ and $\hat{f}(s)$ are the target and predictor networks, respectively.
Reward functions
CAI $r=w_{1}r_{\text{task}}+w_{2}r_{\text{CAI}}+w_{3}r_{\text{reg}}$
RND $r=w_{1}r_{\text{task}}+w_{4}r_{\text{RND}}+w_{3}r_{\text{reg}}$
heuristics $r=w_{1}r_{\text{task}}+w_{5}r_{\text{heu}}+w_{3}r_{\text{reg}}$
$r_{\text{CAI}}$ weight Equation 3
Reward coefficients
Task $w_{1}$ $w_{3}$ $w_{4}$ $w_{5}$ $w_{2,b}$ $\alpha_{1}$ $\alpha_{2}$
Single object 15 -5e-3 10 0.01 40 12e-5 1.5e-6
Single wall 40 -5e-3 10 0.01 40 12e-5 1.5e-6
Multi-wall 40 -5e-3 10 0.01 40 12e-5 1.5e-6
Table C.1: Simulation training reward parameters. All CAI methods (CAI-kinematic, CAI-learned, CAIMAN) use the CAI reward function, all RND methods (RND-full, RND-object) use the RND reward function, while only the Heuristics method uses the heuristics reward function.
Appendix D Additional training details
The low-level policy operates with a control interval of 0.02 seconds, while the high-level policy has a control interval of 0.2 seconds. For every iteration through the high-level control loop, we step the environment by 0.2 seconds, compute the rewards and terminations, collect transitions for the high-level and dynamics policies, and calculate and add the CAI exploration bonus to the rewards buffer. We trained 1500 iterations for the single object and single wall tasks and 2000 iterations for the multi-wall task, where each iteration consists of 10 high-level control steps. All tasks have an episode length of 20 seconds.
In addition to CAI, we also generate meaningful exploration through colored noise during policy rolling out [68, 69]. By sampling from time-correlated actions, we reduce the possibility of meaningless back-and-forth behavior that could result from commonly used white-noise samples. We select a correlation strength parameterized as $\beta=0.5$, corresponding to a colored noise between white and pink.
We list all the hyperparameters used for training the high-level loco-manipulation policy and dynamics residual model in Table D.1.
Environment hyperparameters
number of environments 4096
$v_{r,des}^{x}$ (m/s) $[-1.0,1.0]$
$v_{r,des}^{y}$ (m/s) $[-1.0,1.0]$
$\omega_{r,des}^{z}$ (rad/s) $[-1.0,1.0]$
$\delta{v}_{r}^{x}$ (m/s) $0.3$
$\delta{v}_{r}^{y}$ (m/s) $0.3$
$\delta{\omega}_{r}^{z}$ (rad/s) $0.4$
threshold for kinematic prior model $\epsilon_{p}$ (m) $0.7$
threshold for sparse task reward $\epsilon$ (m) $0.1$
success rate threshold $\epsilon_{s}$ (m) $0.15$
high-level policy frequency (Hz) $5$
low-level policy frequency (Hz) $50$
number of sampled action in CAI $K$ 64
PPO hyperparameters
policy network $[512,256,128]$
policy activation function ELU
policy initial standard deviation $1.0$
value network $[512,256,128]$
value activation function ELU
number of mini-batches $4$
number of epochs per iteration $5$
clip threshold $0.2$
learning rate schedule adaptive
desired KL divergence $0.01$
value function loss coefficient $1.0$
entropy coefficient $0.005$
max gradient norm $1.0$
$\gamma$ $0.99$
$\lambda$ $0.95$
Dynamics hyperparameters
fixed variance $\sigma$ $1.0$
residual model $[256,128]$
batch size $4096$
number of epochs per iteration $8$
learning rate $0.0001$
buffer size $10000000$
Table D.1: High-level policy and dynamics residual training hyperparameters.
Appendix E Domain randomization for hardware deployment
Table E.1 presents the environment parameters that were randomized in simulation during training policies for hardware deployment.
Parameter Range
Object mass $(kg)$ $\mathcal{U}[3.5,7.5]$
Object friction coefficient $\mathcal{U}[0.5,1.5]$
Initial robot position $(m)$ x: $\mathcal{U}[-0.1,0.1]$, y: $\mathcal{U}[-0.1,0.1]$
Initial object position $(m)$ x: $\mathcal{U}[0.9,1.1]$, y: $\mathcal{U}[-0.1,0.1]$
Table E.1: Randomized parameters in simulation for training HL policies for hardware.
Appendix F Additional Experiment: Dense task reward
Though we primarily focused on a sparse task reward, we also trained all tasks in simulation under a dense task reward for all methods. The dense task reward, defined as $r_{\text{task}}^{dense}=\exp{-\left\lVert\bm{p}_{o}-\bm{p}_{t}\right\rVert_{2}}$, awards an amount inversely proportional to the euclidean distance between the object and the target. Combining the dense task rewards with a distance-based heuristic reward resembles the settings from prior work [25].
Task performance during training is shown in Figure F.1. For learning the single object task with dense rewards, all methods demonstrate similar learning speeds and success rates, which indicates that within simpler, less cluttered environments, a dense reward scheme sufficiently guides task learning.
However, for the dense reward single wall and multi-wall tasks, all CAI-imbued learning methods (CAIMAN, CAI-kinematic, CAI-learned) exhibit better sample efficiency over the heuristics method, indicating that learning the actions necessary for obstacle navigation is facilitated through CAI exploration. CAIMAN and CAI-kinematic achieve the fastest convergence and the highest success rates, while CAI-learned is slightly less sample efficient. This can be attributed to the fact that learning dynamics from scratch initially produces erroneous predictions, which results in inaccurately high CAI scores and diverts the robot from acquiring task specific skills. In contrast, with only the kinematic prior model, the resulting exploration behavior is less adversely impacted when skill learning is heavily guided by dense rewards.
HeuristicsCAI-kinematicCAI-learnedCAIMAN
Figure F.1: Policy success rate during training under a dense task reward for all methods. All experiments utilize a minimum of 2 seeds.
Appendix G Additional Experiment: Cart
To demonstrate CAIMAN’s effectiveness in handling complex objects with asymmetric dynamics, we provide an additional experiment: pushing a simplified cart with two normal and two caster wheels, as shown in Figure G.1. The scene configuration is identical to that of the Single object task, with the box being replaced by the cart. Our results demonstrate that CAIMAN is more sample efficient and achieves a higher asymptotic performance (approx. 90% success rate) than most baselines and is on par with the best competitor. Notably, this experiment did not require any changes to the framework or physics prior, highlighting the capability of the learned model in estimating complex dynamics and the effectiveness of CAIMAN for learning with irregular objects. This environment has no obstacles, but we hypothesize that cluttered settings with walls would further emphasize CAIMAN’s advantages over baselines.
HeuristicsCAI-kinematicCAI-learnedRND-fullRND-objectCAIMAN
Figure G.1: Left: A simplified cart object. Right: Single cart pushing results.

### v3

@@
 Enabling legged robots to perform non-prehensile loco-manipulation is crucial for enhancing their versatility.
-Learning behaviors such as whole-body object pushing often requires sophisticated planning strategies or extensive task-specific reward shaping, especially in unstructured environments.
+However, learning behaviors such as whole-body object pushing often necessitates sophisticated planning strategies or extensive task-specific reward shaping.
 In this work, we present CAIMAN, a practical reinforcement learning framework that encourages the agent to gain control over other entities in the environment.
@@
 We empirically demonstrate CAIMAN’s superior sample efficiency and adaptability to diverse scenarios in simulation, as well as its successful transfer to real-world systems without further fine-tuning.
-Keywords: Reinforcement Learning, Loco-manipulation, Intrinsic Motivation
-1 Introduction
-Modern legged robots continue to impress with their diverse capabilities, from traversing challenging terrains [1, 2, 3] to executing agile maneuvers such as backflips and parkour [4, 5, 6].
-However, enabling these systems to interact with their environments remains an open research challenge [7, 8, 9].
-A common approach is to equip legged robots with external manipulators for prehensile manipulation [10, 11, 12, 13].
-Although effective, these methods are often constrained by the size and weight of objects they can successfully manipulate.
-In contrast, non-prehensile manipulation demands whole-body coordination and remains a non-trivial task.
-Traditional approaches for whole-body manipulation explicitly model robot-object interactions, often relying on complex planning and optimization for coordinating motion [14].
-These methods require accurate models of both the robot and the environment, making them less scalable to high-dimensional systems with complex, stochastic dynamics, and often limiting contact to predefined regions [15, 16].
-Learning-based methods offer a more scalable alternative, improving computational efficiency and reducing dependence on precise object estimation [17, 18, 19].
+A video demo is available at https://www.youtube.com/watch?v=dNyvT04Cqaw.
+I INTRODUCTION
+Modern day legged robots showcase impressive versatility, from traversing challenging terrains [miki2022learning, choi2023learning] to executing agile maneuvers such as backflips and parkour [katz2019mini, cheng2024extreme].
+Yet as expectations for autonomy grow, enabling these systems to physically interact with their environments remains a central research problem [ha2024learning, tang2024deep, gu2025humanoid].
+A common strategy to enhance manipulation capabilities is to equip legged robots with external manipulators for prehensile tasks [bellicoso2019alma, zimmermann2021go, sleiman2021unified], but such methods are inherently constrained by object size and payload.
+Harnessing whole-body motion for non-prehensile manipulation offers a promising alternative, but remains a non-trivial challenge.
+Traditional approaches for whole-body manipulation explicitly model robot-object interactions, often relying on complex planning and optimization for coordinating motion [murooka2015whole].
+These methods require accurate models of both the robot and the environment, which restricts their scalability for high-dimensional systems with complex, stochastic dynamics and limits contact points to predefined regions [farnioli2016toward, rigo2023contact].
+Learning-based methods offer a more scalable alternative, improving computational efficiency and reducing dependence on precise object estimation [shi2021circus, cheng2023legs, arm2024pedipulate].
 However, these methods frequently rely on tedious reward shaping or learning curriculums to guide exploration and foster meaningful behavior.
-Furthermore, the large exploration space of loco-manipulation tasks often necessitates special treatment, such as learning from behavioral priors  [20, 21, 22] or task-agnostic explorative rewards [23, 24], to encourage meaningful engagement with the environment.
-Prior works have also leveraged a hierarchical structure to decouple locomotion and high-level planning [25].
-Training the high-level policy, however, still requires elaborate reward design and shaping.
+Furthermore, the large exploration space of loco-manipulation tasks often necessitates special treatment, such as learning from behavioral priors  [ha2024umi, urpi2023efficient] or task-agnostic explorative rewards [schwarke2023curiosity, zhang2024wococo], to encourage meaningful engagement with the environment.
+Figure 1: Robot loco-manipulation in the real world enabled by CAIMAN. The quadruped maneuvers a box around an obstacle to reach a target position.
 To alleviate these challenges, we introduce CAIMAN, a framework for training non-prehensile object pushing skills in legged robots.
@@
 We train robust locomotion using an existing pipeline, and we learn the high-level policy with a simple yet effective reward structure consisting of only three terms: a sparse task reward, an action regularizer, and an intrinsically motivated explorative reward.
-The sparse task reward provides a general signal for various pushing tasks, including ones requiring navigation through obstacles.
-The high-level policy is incentivized to explore and gain control over the environment via an exploratory reward based on Causal Action Influence (CAI) [26], which quantifies how much influence the agent has on the states of other entities and is computed based on environment dynamics.
-Instead of learning the dynamics from scratch, CAIMAN utilizes a simple pre-defined kinematic prior, and combines it with learned residual dynamics that compensate for physical interactions beyond the prior model.
-Using only the sparse task reward, we show that CAIMAN achieves strong learning performance and superior sample efficiency against other baselines, especially in complex scenarios with obstacles.
+The sparse task reward provides a general signal for various pushing tasks, where the robot is only rewarded for completing the task successfully.
+The exploratory reward incentivizes the robot to explore and gain control over the environment via Causal Action Influence (CAI) [seitzer2021causal], which is a measure quantifying how much influence an agent has on the states of other entities and is computed based on environment dynamics.
+Instead of learning the dynamics from scratch, CAIMAN combines a simple predefined kinematic prior with learned residual dynamics that compensate for physical interactions beyond the prior, allowing accurate dynamics to be learned efficiently.
+We show that in a sparse task reward setting, CAIMAN achieves strong learning performance and superior sample efficiency compared to other baselines, including within complex, obstacle-laden scenarios.
 We also demonstrate that the learned residual accurately captures complex robot-object interactions that are not modeled by the kinematic prior.
-Finally, we successfully transfer the learned policy to a real quadruped robot, enabling it to perform whole-body object pushing in obstacle environments.
-To summarize, our contribution is three-fold: 1) a general hierarchical framework for learning whole-body object pushing with legged robots in various scenarios, including navigating through obstacles; 2) an intrinsically motivated reward based on causal influence calculated from a combination of a kinematics prior and learned residual dynamics; 3) Successful hardware validation on a real-world quadruped robot.
-2 Related Work
-2.1 Loco-manipulation on Legged Systems
+Finally, we successfully transfer the learned policy to a real quadruped robot, enabling it to perform whole-body object pushing in real-world settings.
+To summarize, our contribution is three-fold: 1) a general hierarchical framework for learning whole-body object pushing with legged robots in various scenarios, including navigation through obstacles; 2) an intrinsically motivated reward based on causal influence calculated from a combination of a kinematics prior and learned residual dynamics; 3) successful hardware validation on a real-world quadruped robot.
+II RELATED WORK
+II-A Loco-manipulation on Legged Systems
 The integration of locomotion and manipulation has gained significant attention as a promising and application-oriented research field for legged robots.
-Traditional model-based methods that rely on accurate representations of both robot and object to optimize trajectories [11, 14, 16, 27] have shown success in tasks such as box carrying [10, 12], but often require accurate state estimation [28, 29, 30] and struggle with scalability due to the complexity of contact modes [31, 32].
-Reinforcement learning (RL) is a promising alternative that avoids explicit modeling and has been successfully applied to tasks such as soccer dribbling [33, 34], button pushing [18, 35], and door opening [19, 23, 36].
-However, effective exploration may be difficult to achieve due to the large search space of robot-object interactions [17, 37].
-To address exploration challenges, researchers have proposed using behavior prior [38, 39] or task-agnostic explorative rewards [23, 24] as mechanisms to guide learning.
-Furthermore, hierarchical frameworks [40, 41] are utilized to decompose tasks into high-level planning and low-level control, enabling effective behavior across various scenarios [42, 43].
+Model-based methods that rely on accurate representations of both robot and object to optimize trajectories [zimmermann2021go, murooka2015whole, rigo2023contact] have shown success in tasks such as box carrying [bellicoso2019alma, sleiman2021unified], but often require precise state estimation [lin2024locoman, schakkal2024dynamic] and struggle with scalability due to the complexity of contact modes [cheng2023enhancing].
+Reinforcement learning (RL) is a viable alternative that avoids explicit modeling and has been successfully applied to diverse tasks including soccer dribbling [ji2023dribblebot, hu2024dexdribbler], button pushing [cheng2023legs, he2024learning], and door opening [arm2024pedipulate, schwarke2023curiosity].
+However, achieving effective exploration remains difficult given the large search space of robot-object interactions [shi2021circus, fu2023deep].
+To address exploration challenges, researchers have proposed using behavior prior [sleiman2024guided, liu2024opt2skill] or task-agnostic explorative rewards [schwarke2023curiosity, zhang2024wococo] as mechanisms to guide learning.
+In addition, hierarchical frameworks [rigo2024hierarchical, wang2024hypermotion] decompose tasks into high-level planning and low-level control, enabling effective behaviors across various scenarios [kumar2023cascaded].
 In accordance with previous works, we also implement a hierarchical control framework for whole-body object pushing, decoupling high-level velocity planning from low-level locomotion, similarly to Jeon et al.
-[25].
-We aim to utilize a simple reward structure, as opposed to one that requires sophisticated design effort as seen in [25], to achieve sample-efficient high-level policy training, even under complex scenarios such as object manipulation through obstacles.
-2.2 Intrinsically Motivated Reinforcement Learning
-Intrinsic motivation (IM) [44] plays a vital role in reinforcement learning, especially when extrinsic rewards are sparse or difficult to design.
-Core IM mechanisms include curiosity [45], learning progress [46], and empowerment [47], each promoting exploration and skill acquisition.
-Curiosity, often measured via prediction errors in learned world models [48, 49, 50], rewards agents for encountering novel states, and has been previously applied for learning loco-manipulation skills [23, 24].
-Learning progress encourages agents to focus on regions in the state space with rapid improvement, supporting curriculum learning and adaptive exploration [51, 52].
-Our work builds on empowerment, an information-theoretic quantity defined as the channel capacity between the agent’s actions and its sensory observations [47, 53, 54].
-Recent work has connected IM to causal reasoning, aiming to improve sample efficiency and interpretability [55, 56, 57, 58, 59].
-To encourage effective exploration, we employ causal action influence (CAI) [26], a measure of an agent’s ability to influence its environment and a conceptual lower bound on empowerment.
-3 Preliminaries
-(a) No influence of $A$ on $S^{\prime}_{2}$
-(b) Influence of $A$ on $S^{\prime}_{1}$ and $S^{\prime}_{2}$
+[jeon2023learning].
+We aim to utilize a simple reward structure, as opposed to one that requires sophisticated design effort as seen in [jeon2023learning], to achieve sample-efficient high-level policy training.
+II-B Intrinsically Motivated Reinforcement Learning
+Intrinsic motivation (IM) [rm2000intrinsic] plays a vital role in reinforcement learning, especially when extrinsic rewards are sparse or difficult to design.
+Core IM mechanisms include curiosity [schmidhuber1991possibility], learning progress [schmidhuber2010formal], and empowerment [klyubin2005empowerment], each promoting exploration and skill acquisition.
+Curiosity, often measured via prediction errors in learned world models [pathak2017curiosity, burda2019exploration], rewards agents for encountering novel states, and has been previously applied for learning loco-manipulation skills [schwarke2023curiosity, zhang2024wococo].
+Learning progress encourages agents to focus on regions in the state space with rapid improvement, supporting curriculum learning and adaptive exploration [blaes2019control, colas2019curious].
+Our work builds on empowerment, an information-theoretic quantity defined as the channel capacity between the agent’s actions and its sensory observations [klyubin2005empowerment, mohamed2015variational].
+Recent work has connected IM to causal reasoning, aiming to improve sample efficiency and interpretability [buesing2018woulda, sontakke2021causal].
+To encourage effective exploration, we employ causal action influence (CAI) [seitzer2021causal], a measure of an agent’s ability to influence its environment and a conceptual lower bound on empowerment.
+III PRELIMINARIES
+(a) No influence of AA on $S^{\prime}_{2}$
+(b) Influence of AA on $S^{\prime}_{1}$ and $S^{\prime}_{2}$
 Figure 2: Illustration of the LCM (left) for two different environment situations $S=s$ (right) in the loco-manipulation task.
 The LCM captures the transition from $S,A$ to $S^{\prime}$, factorized into state components.
-While the global SCM is fully connected (dashed and continuous lines), the LCM $\mathcal{G}_{S=s}$ (continuous lines) is causally minimal.
+The global SCM is fully connected (dashed and continuous lines), while the LCM $\mathcal{G}_{S=s}$ (continuous lines) is sparser.
 We are interested in detecting the presence of continuous orange arrows in the LCM, i.e.
-the influence of the action $A$ on next states $S^{\prime}$.
-We model decision-making in a dynamic environment as a Markov Decision Process (MDP) [60], defined by the tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $P$ the transition kernel, $R$ the reward function, and $\gamma$ the discount factor.
-Following the principle of independent causal mechanisms [61], we assume that the world consists of interacting but independent entities.
+the influence of the action AA on next states $S^{\prime}$.
+We model decision-making in a dynamic environment as a Markov Decision Process (MDP) [sutton2018reinforcement], defined by the tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $P$ the transition kernel, $R$ the reward function, and $\gamma$ the discount factor.
+Following the principle of independent causal mechanisms [peters2017elements], we assume that the world consists of interacting but independent entities.
 This induces a state space factorization $\mathcal{S}=\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{N}$ for $N$ entities, where each factor $\mathcal{S}_{i}$ represents the state of entity $i$.
-An MDP coupled with a policy $\pi:\mathcal{S}\mapsto\mathcal{A}$ induces a Structural Causal Model (SCM) [62] describing the resulting trajectory distribution.
-Definition 1 (Structural Causal Model [62]).
-An SCM is a tuple $(\mathcal{U},\mathcal{V},F,P^{u})$, where $\mathcal{U}$ are set of exogenous variables (e.g., latent randomness) sampled from $P^{u}$, $\mathcal{V}$ are set of observed variables (e.g., states, actions, rewards), $F$ is the set of structural functions capturing the causal relations, such that functions $f_{V}:\mathrm{Pa}(V)\times\mathcal{U}\rightarrow V$, with $\mathrm{Pa}(V)\subset\mathcal{V}$ denoting the set of parents of $V$, determine the value of endogenous variables $V$ for each $V\in\mathcal{V}$.
+An MDP coupled with a policy $\pi:\mathcal{S}\mapsto\mathcal{A}$ induces a Structural Causal Model (SCM) [pearl2009causality] describing the resulting trajectory distribution.
+Figure 3: CAIMAN framework: The high-level (HL) policy generates desired base velocity commands, which are translated into joint commands by a low-level ($L$) policy.
+We utilize a simple kinematic prior and learned residual dynamics to model the robot-object interaction in the environment while providing a CAI-based explorative bonus along with the sparse task reward.
+Definition 1 (Structural Causal Model [pearl2009causality])
+An SCM is a tuple $(\mathcal{U},\mathcal{V},F,P^{u})$, where $\mathcal{U}$ is a set of exogenous variables (e.g., latent randomness) sampled from $P^{u}$, $\mathcal{V}$ is a set of observed variables (e.g., states, actions, rewards), and $F$ is the set of structural functions capturing the causal relations, such that functions $f_{V}:\mathrm{Pa}(V)\times\mathcal{U}\rightarrow V$, with $\mathrm{Pa}(V)\subset\mathcal{V}$ denoting the set of parents of $V$, determine the value of endogenous variables $V$ for each $V\in\mathcal{V}$.
 SCMs are typically visualized as directed acyclic graphs, where nodes represent variables and edges indicate causal relations, as shown in Figure 2.
-In our case, the SCM captures one-step transitions with variables $\mathcal{V}={S_{1},\ldots,S_{N},A,S_{1}^{\prime},\ldots,S_{N}^{\prime}}$.
+The SCM we consider captures one-step transitions with variables $\mathcal{V}={S_{1},\ldots,S_{N},A,S_{1}^{\prime},\ldots,S_{N}^{\prime}}$.
 Due to the Markov property and flow of time, causal dependencies exist only from $(S,A)$ to $S^{\prime}$.
-While most entity pairs may theoretically interact (i.e., $S_{i}/A\rightarrow S^{\prime}_{j}$ for most $i,j$), interactions often become sparse when we observe a specific state configuration.
-We use the Local Causal Model (LCM) to capture such context-specific structure [63].
-Definition 2 (Local Causal Model [63]).
+The global SCM is generally fully connected, encoding all possible, however unlikely, interactions between entities (i.e., $S_{i}/A\rightarrow S^{\prime}_{j}$ for many $i,j$) (see  Figure 2).
+However, most of these dependencies are inactive given a specific state configuration.
+For example, the robot can only influence an object if it is within a certain proximity to it.
+To capture such context-specific structure, we use the Local Causal Model (LCM) [pitis2020counterfactual].
+Definition 2 (Local Causal Model [pitis2020counterfactual])
 Given an SCM $(\mathcal{U},\mathcal{V},F,P^{u})$ and an observation $V=v,V\subset\mathcal{V}$, the local SCM is the SCM with $F_{V=v}$ and graph $\mathcal{G}_{do(V=v)}$ obtained by pruning edges from $\mathcal{G}_{do(V=v)}$ until it is causally minimal.
-In this work, we build on the insight that encouraging agents to gain control over their environment facilitates learning in loco-manipulation tasks.
-We do so by driving the agent toward states where it can influence other entities.
-We make use of a principled explicit measure of local influence, the Causal Action Influence (CAI) [26, 64] measure.
-CAI is a state-dependent measure of control that assesses whether an agent, through its actions, can affect other entities.
-Graphically, it corresponds to predicting the existence of an edge $A\rightarrow S^{\prime}_{j}$ in the LCM $\mathcal{G}_{do(S=s)}$.
-This influence is measured with point-wise conditional mutual information $I(S_{j}^{\prime};A\mid S=s)$.
-4 Method
-Figure 2: CAIMAN framework: The high-level (HL) policy generates desired base velocity commands, which are translated into joint commands by a low-level (LL) policy.
-We utilize a simple kinematic prior and learned residual dynamics to model the robot-object interaction in the environment while providing a CAI-based explorative bonus along with the sparse task reward.
-The hierarchical framework adopted by CAIMAN is presented in Figure 2.
+Intuitively, the LCM is sparser than the global SCM, as it retains only the edges corresponding to entity influences that are active in the current context.
+This ensures that the graph reflects the agent’s local influences, rather than all potential ones.
+In this work, we build on the insight that agents can effectively learn loco-manipulation tasks when encouraged to gain control over their environment.
+We do so by driving the agent toward states where it can influence other entities, achieved through a principled explicit measure of local influence, the Causal Action Influence (CAI) [seitzer2021causal, urpicausal] measure.
+CAI is a state-dependent measure of control that assesses whether an agent’s actions can affect the states of other entities, corresponding graphically to the presence of an edge $A\rightarrow S^{\prime}_{j}$ in the LCM $\mathcal{G}_{do(S=s)}$.
+Formally, CAI is defined via point-wise conditional mutual information $I(S_{j}^{\prime};A\mid S=s)$, which equals zero when $S_{j}^{\prime}$ is independent of $A$ given $S=s$, i.e.
+$S_{j}^{\prime}\perp\!\!\!\perp A\mid S=s$.
+At state $S=s$, CAI is given by
+$\begin{split}C^{j}(s)&:=I(S^{\prime}_{j};A\mid S=s)\\ &=\mathbb{E}_{a\sim\pi}\big[D_{KL}\big(P_{S^{\prime}_{j}\mid S=s,A=a}\;\big|\big|\;P_{S^{\prime}_{j}\mid S=s}\big)\big].\end{split}$ (1)
+IV METHOD
+The hierarchical control and training framework adopted by CAIMAN is presented in Figure 3.
 It consists of a high-level (HL) policy that generates desired base velocity commands and a low-level policy that translates velocity commands into joint-level actions.
-In this work, we introduce a novel, sample-efficient approach for learning the high-level policy, while building upon an existing pipeline [65] to train robust low-level policies.
-4.1 Low-level Locomotion Policy
+In this work, we introduce a novel, sample-efficient approach for learning the high-level policy, while building upon an existing pipeline [rudin2022learning] to train robust low-level policies.
+IV-A Low-level Locomotion Policy
 The low-level locomotion policy $\bm{\pi}_{l}$ is trained to generate target joint positions $\bm{q}_{j,des}\in\mathbb{R}^{12}$ that track a desired base velocity command $\bm{v}_{r,des}\in\mathbb{R}^{3}$.
@@
 The policy is trained with velocity commands sampled uniformly from a predefined range.
-To ensure robust sim-to-real transfer and hardware performance during contact-rich tasks like pushing, we apply domain randomization and external perturbations following [65].
-Further details of the low-level observation $\bm{o}_{l}$ can be found in Appendix A.
-4.2 High-level Planning Policy
-The high-level policy $\bm{\pi}_{h}$ is trained to generate desired robot velocity commands $\bm{a}_{h}:=\bm{v}_{r,des}$ that achieve successful task completion in object pushing.
-Its observation $\bm{o}_{h}$ includes the robot’s linear and angular velocities, the poses of the robot and object, the target object position $\bm{p}_{t}=(x_{p},y_{p})\in\mathbb{R}^{2}$, and the previous action—all expressed in the world frame.
-In scenarios with obstacles (e.g., walls), $\bm{o}_{h}$ also includes their poses in the world frame.
-For a detailed description of the high-level observation, refer to Appendix A.
-We use Proximal Policy Optimization (PPO) [66] as the base RL algorithm, with a simple reward function composed of three terms:
-$r=w_{1}\mathds{1}_{\left\lVert\bm{p}_{o}-\bm{p}_{t}\right\rVert_{2}<\epsilon}+w_{2}r_{\text{CAI}}+w_{3}\left\lVert\bm{a}_{h}-\bm{a}_{h,prev}\right\rVert_{2}^{2},$ (1)
-where $\bm{p}_{o}$, $\bm{p}_{t}$ are the current and target object positions, $\epsilon$ is a success threshold, and $\bm{a}_{h}$, $\bm{a}_{h,prev}$ are the current and previous high-level actions, respectively.
-The exploration bonus $r_{\text{CAI}}$ is derived from the CAI measure $\tilde{C}_{j}$ over the object $j$.
-We define the CAI reward as:
-$r_{\text{CAI}}=\tilde{C}_{j=\text{object}}(s)=\frac{1}{K}\sum_{i=1}^{K}D_{\mathrm{KL}}\left(P_{S_{j}^{\prime}\mid S=s,A=a^{(i)}}\bigg{\|}\frac{1}{K}\sum_{k=1}^{K}P_{S_{j}^{\prime}\mid S=s,A=a^{(k)}}\right),$ (2)
-given $K$ actions $\{a^{(i)}\}_{i=1}^{K}$ sampled from the policy.
-This approximation estimates the marginal $P_{S_{j}^{\prime}\mid s}$ using Monte Carlo sampling.
-We model the transition distribution $P_{S_{j}^{\prime}\mid S=s,A=a}$ as a fixed-variance Gaussian (details in Section 4.3), which enables closed-form KL divergence calculation using the Gaussian mixture approximation [67].
-The CAI reward encourages the agent to reach states where it exerts greater influence over the object, thereby promoting task-relevant exploration and learning.
-The weight $w_{2}$ for the CAI reward scales with the raw CAI score:
-$w_{2}=w_{2,b}+\max\left(0,(r_{\text{CAI}}-\alpha_{1})/\alpha_{2}\right),$ (3)
-where $\alpha_{1}$ is the threshold for scaling and $\alpha_{2}$ controls the scaling rate.
-Finally, we foster more directed exploration by injecting time-correlated noise into the action sampling process during training [68, 69].
-For more details on training the high-level policy, we refer the reader to Appendix D.
-4.3 Dynamics learning
-To calculate the exploration bonus $r_{\text{CAI}}$ in (2), we learn the transition model $P_{S_{j}^{\prime}\mid S=s,A=a^{(k)}}$ for all entities $j$.
-In our current setting, we focus on a single entity—the pushable object—but the framework naturally extends to multiple entities.
-Concretely, this reduces to learning the object transition model $P_{\bm{s}^{\prime}_{o}\mid\bm{s}_{h},\bm{a}_{h}}$.
-The object state $\bm{s}_{o}$ is defined as the object’s position, while the high-level state $\bm{s}_{h}=(\bm{\xi}_{r},\bm{\xi}_{o},\bm{v}_{r},\bm{v}_{o})$ includes the robot’s pose $\bm{\xi}_{r}$ and velocity $\bm{v}_{r}$ and the object’s pose $\bm{\xi}_{o}$ and velocity $\bm{v}_{o}$.
-The high-level action $\bm{a}_{h}$ corresponds to the desired robot velocity, and the next object state $\bm{s}^{\prime}_{o}=\bm{p}^{\prime}_{o}=(x^{\prime}_{o},y^{\prime}_{o})$ denotes its position at the subsequent timestep.
-Instead of using a pretrained model, CAIMAN leverages data collected from high-level interactions during training to efficiently learn the dynamics.
-As described earlier, we model the object’s transition probability $P_{\bm{s}^{\prime}_{o}\mid\bm{s}_{h},\bm{a}_{h}}$ as a fixed-variance Gaussian distribution $\mathcal{N}(\bm{s^{\prime}}_{o};f_{\theta}(\bm{s}_{h},\bm{a}_{h}),\sigma^{2})$, where $f_{\theta}$ is a neural network that predicts the mean of the distribution.
-To enhance learning efficiency, we incorporate a simple kinematic prior model $f_{\text{kin}}$, which estimates the next object position $\tilde{\bm{p}}^{\prime}_{o}$ using geometric reasoning based on the relative pose between robot and object and the commanded velocity.
-This estimate is computed by projecting the robot’s velocity $\bm{a}_{h}$ onto the direction from the robot to the object and updating the object’s position accordingly:
-$\tilde{\bm{p}}^{\prime}_{o}=\begin{cases}\bm{p}_{o}+\delta t\cdot(\bm{a}_{h,xy}\cdot\delta\hat{\bm{p}})\cdot\delta\hat{\bm{p}},\quad&\text{if }\left\lVert\bm{p}_{o}-\bm{p}_{r}\right\rVert_{2}\leq\epsilon_{p}\text{ and }\bm{a}_{h,xy}\cdot\delta\hat{\bm{p}}>0\\ \bm{p}_{o},&\text{otherwise.}\end{cases}$ (4)
-Here, $\delta\hat{\bm{p}}=(\bm{p}_{o}-\bm{p}_{r})/\left\lVert\bm{p}_{o}-\bm{p}_{r}\right\rVert_{2}$ is the unit vector pointing from the robot to the object, $\delta t$ is the high-level control step size, and $\epsilon_{p}$ is a distance threshold.
-The conditions ensure that the robot is close enough to and moving toward the object.
-We combine the kinematic prior with a learned residual model $f_{\text{res}}$, parameterized by $\theta$, to capture complex physical interactions beyond the kinematics model, including nonlinear effects such as friction, drag, and collisions with obstacles.
-The final dynamics model is thus:
-$f_{\theta}(\bm{s}_{h},\bm{a}_{h})=f_{\text{kin}}(\bm{s}_{h},\bm{a}_{h})+f_{\text{res}}(\bm{s}_{h},\bm{a}_{h};\theta),$ (5)
-where $f_{\text{kin}}$ is deterministic and independent of $\theta$. We train the residual model by minimizing the mean squared error between the predicted and true object positions:
-$\mathcal{L}_{\text{res}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\lVert\tilde{\bm{s}}^{\prime(i)}_{o}-\bm{s}^{\prime(i)}_{o}\right\rVert_{2}^{2},$ (6)
-where $\tilde{\bm{s}}^{\prime(i)}_{o}=f_{\text{kin}}(\bm{s}^{(i)}_{h},\bm{a}^{(i)}_{h})+f_{\text{res}}(\bm{s}^{(i)}_{h},\bm{a}^{(i)}_{h};\theta)$.
-Although the high-level action $\bm{a}_{h}$ could, in principle, be sampled from the full support of the desired velocity $\bm{v}_{r,des}$ [26], not all commands are feasible for the low-level controller to execute under the robot’s current state.
-Therefore, we define $\bm{a}_{h}=\tilde{\bm{v}}_{r}$ to be the achievable velocity, and use $\tilde{\bm{v}}_{r}$ for both dynamics learning and CAI computation.
-Given training samples $\mathcal{D}={(\bm{s}_{h}^{(i)},\bm{a}_{h}^{(i)},\bm{s}^{\prime(i)}_{h})}$, we have access to the achieved robot velocity $\bm{v}^{\prime}_{r}$ at the next timestep.
-When computing CAI as an exploration reward, we assume that the robot’s velocity can only change within a limited range over one high-level step.
-Thus, for any state $\bm{s}_{h}$, we sample the action $\bm{a}_{h}$ from a bounded range centered on the current velocity:
-$\bm{a}_{h}=\tilde{\bm{v}}_{r}\sim\mathcal{U}[\bm{v}_{r}\pm\delta\bm{v}_{r}]$ (7)
-where $\delta\bm{v}_{r}$ is a fixed velocity deviation range. Detailed values for the velocity limits $(\delta v_{r}^{x},\delta v_{r}^{y},\delta\omega_{r}^{z})$ are provided in Appendix D.
-5 Experiments
-Figure 3: Illustrations for the Single-object (left), Single-wall (middle), and Multi-wall (right) tasks. The yellow sphere denotes the object’s target position.
-We train both the high-level and low-level policies using Isaac Sim [70].
-Training hyperparameters and additional details are provided in Appendix D.
-We evaluate CAIMAN on three pushing tasks of increasing difficulty, illustrated in Figure 3.
-For the Single-object task, the robot must manipulate a single cuboid object to a fixed target position.
-The Single-wall task includes a fixed wall that blocks the direct path between the object and the target.
-The robot must navigate the object around the wall to reach the goal.
-Finally, the Multi-wall task includes multiple fixed walls that obstruct the path between the object and the target.
-The robot must maneuver the object through these obstacles to complete the task.
-The presence of fixed wall obstacles in the single-wall and multi-wall tasks introduces regions in the state space where object dynamics become more complex, reducing the robot’s ability to exert consistent influence over the object.
-For all tasks, we train with a fixed target position.
-A full description of scene configurations is provided in Appendix B.
-Baselines
-We compare our approach against several baselines.
-The Heuristics baseline trains the high-level policy using a distance-based heuristic reward, $r_{\text{heu}}=\exp\left(-\left\lVert\bm{p}_{r}-\bm{p}_{o}\right\rVert_{2}\right)$, encouraging the robot to minimize its distance to the object.
-This mirrors the main explorative reward from prior work [25].
-We also conduct comparisons with other intrinsic motivation algorithms, namely Random Network Distillation (RND) [71].
-We implement two variants: RND-full, which mirrors the original RND implementation and computes the prediction error for the full state, and RND-object, which computes the prediction error for the object’s state only.
-As for ablations, we implement CAI-kinematic, where CAI is computed using only the kinematic prior (Equation 4), and and CAI-learned, where CAI is computed using a fully learned dynamics model (no prior).
-For detailed formulations of all reward functions and their coefficients, see Appendix C.
-5.1 Simulation results
-HeuristicsCAI-kinematicCAI-learnedRND-fullRND-objectCAIMAN
-Figure 4: Policy success rate evaluated at every 100 training iterations for all methods and tasks.
-Results are evaluated across 800 episodes and averaged over 3 seeds, shaded area represents standard deviation.
-We evaluate the learning performance of each method by measuring the success rate for each task.
-A task is considered successfully completed if the distance between the object and the target is below a threshold $\epsilon_{s}$ at the end of an episode.
-The results are shown in Figure 4.
-With only sparse task rewards, task-relevant skill acquisition heavily relies on effective exploration.
-In the Single-object task, all methods—except RND-full—achieve some degree of success.
-CAIMAN, RND-object, and Heuristics perform best, while CAI-kinematic and CAI-learned underperform.
-The low performance of CAI-learned is likely due to the high data and time demands required to learn an accurate dynamics model from scratch.
-In contrast, CAI-kinematic performs comparably to the top methods, suggesting that the naive kinematic prior is sufficient for simpler tasks with straightforward object interactions.
-In the more complex Single-wall and Multi-wall tasks, only CAIMAN and RND-object demonstrate notable success, with RND-object exhibiting lower asymptotic performance than CAIMAN in both scenarios.
-This highlights CAIMAN’s unique ability to guide task-relevant exploration even in the presence of obstacles, validating its effectiveness in cluttered environments without relying on dense extrinsic rewards.
-The failure of CAI-kinematic and CAI-learned in these tasks emphasizes the importance of combining a kinematic prior with a learned residual model to efficiently capture the complex object dynamics needed for effective exploration.
-RND-full results in agents randomly exploring the state space without yielding meaningful behavior in all tasks.
-This finding reinforces the need for object-centric exploration bonuses when the skill to be learned involves physical interaction with objects.
-The performance gap between RND-object and CAIMAN can be partially attributed to the phenomenon of detachment [72], where the agent drifts away from reward-depleted areas and continuously pursues new regions with higher intrinsic reward, potentially losing avenues of task-relevant exploration.
-5.2 Leveraging pretrained dynamics for new tasks
-Figure 5: Single wall task with target positions randomly sampled within an area in front of the wall.
-Figure 6: Learning the single wall random target task using learning from scratch and learning with models transferred from the fixed target task.
-Since our framework learns the environment dynamics, we propose that the dynamics residual obtained from a previous training can be reused across different tasks—as long as the underlying dynamics remain consistent.
-We hypothesize that leveraging a pretrained model improves the accuracy of CAI estimates in the early stages of training, thereby enhancing the exploration signal and significantly boosting sample efficiency.
-To test this hypothesis, we consider a generalization of the Single-wall task in which the target position is randomized instead of fixed, sampled uniformly within a predefined area as illustrated in Figure 6.
-Detailed scene configurations are provided in Appendix B.
-We compare the original CAIMAN, which learns the dynamics residual from scratch, to one that reuses a pretrained residual model.
-For the latter, we initialize the residual with one from the original Single-wall task and continuously fine-tune it during training with data collected from the new generalized task.
-As shown in Figure 6, the reuse of learned dynamics significantly accelerates learning with sparse rewards.
-The pretrained model yields more informative CAI estimates, guiding the robot toward meaningful interaction behaviors early in training.
-This leads to substantial gains in sample efficiency, supporting our hypothesis and demonstrating the benefits of transferring learned dynamics across related tasks.
-5.3 Hardware deployment
-Figure 7: Snapshots from the hardware deployment for the single object and single wall tasks. The mass of the box is 5.5 kilograms with a dimension of $(0.55,0.55,0.5)$ meters.
-We validated our trained policies on real-world quadruped pushing tasks with Unitree Go2, as shown in Figure 7.
-To bridge the sim-to-real gap, we applied domain randomization to object mass and friction during high-level policy training.
-Additional details are provided in Appendix E.
-All entities in the environment were tracked using an external motion capture system.
-The trained policies are deployed directly to the robot and successfully execute pushing tasks without requiring additional fine-tuning.
-We refer interested readers to the supplementary video for more details.
-6 Conclusion
-We present CAIMAN, a general framework for training whole-body pushing skills in legged robots.
-CAIMAN adopts a hierarchical control strategy that decouples locomotion from high-level planning.
-By introducing an intrinsic exploration bonus based on CAI, our method encourages the robot to gain control over relevant entities in its environment under sparse task signals—eliminating the need for tedious, hand-crafted reward shaping, even in cluttered scenarios with obstacles.
-To further enhance training efficiency, we bootstrap the CAI computation using a simple yet effective kinematic prior, and refine it with a learned residual dynamics model.
-We evaluate CAIMAN on a suite of quadruped pushing tasks and show that it consistently outperforms competitive baselines in terms of sample efficiency.
-We also demonstrate that the learned dynamics residual can be reused to accelerate training for new tasks, highlighting the transferability of our approach.
-Finally, we validate our method on real hardware, seamlessly deploying the trained policies without the need for fine-tuning.
-Overall, CAIMAN offers a robust and scalable solution for learning physically grounded manipulation behaviors in legged robots, paving the way toward more autonomous and adaptable robotic systems.
-7 Limitations
-Although CAI-based exploration provides meaningful guidance for robot-object interaction, it inevitably relies heavily on an accurate dynamics model.
-Currently, CAIMAN trains the high-level policy alongside the dynamics residual, which may lead to undesired exploratory behaviors in the early stages of training when the dynamics model is still inaccurate.
-This can be mitigated by leveraging a pretrained world model.
-Our deployment pipeline currently relies on an external motion capture system.
-We leave the incorporation of an onboard perceptive module for future work.
-Another potential direction for the future is to extend CAIMAN to enable exploration for multi-object interactions.
-Acknowledgments
-We thank Dongho Kang, Taerim Yoon, Zijun Hui, Flavio De Vincenti and Hélène Stefanelli for developing the hardware testing framework.
-Appendix A Detailed Observation Space
-Below we list the observation space for low-level, high-level policies and the high-level state for calculating CAI.
+To ensure robust sim-to-real transfer and hardware performance during contact-rich tasks like pushing, we apply domain randomization and external perturbations following [rudin2022learning].
+The complete low-level observation space $\bm{o}_{l}$ is shown in Table I.
 Low-level observation $\bm{o}_{l}$
@@
 $_{\mathcal{B}}\bm{g}\in\mathbb{R}^{3}$ Projected gravity in base frame $\mathcal{B}$
-$\bm{v}_{r,des}=(v^{x}_{r,des},v^{y}_{r,des},\omega^{z}_{r,des})\in\mathbb{R}^{3}$ Desired velocity command
+$\bm{v}_{r,des}\in\mathbb{R}^{3}$ Desired velocity command
 $\bm{a}_{l,prev}\in\mathbb{R}^{12}$ Previous action
@@
 $\bm{v}_{o}=(v_{o}^{x},v_{o}^{y},v_{o}^{z})\in\mathbb{R}^{3}$ Object velocity
-Table A.1: Detailed observation space for each module.
-Appendix B Scene Configurations
-The detailed initial pose for each entity and target position are shown in Table B.1.
-In simulation, the object measures $0.5$ meters in length, width, and height, and weighs $5.0$ kilograms.
-The wall in the single-wall task has dimensions $(0.1,1.0,0.5)$ meters, while walls in the multi-wall task measure $(0.1,1.25,0.5)$ meters.
-For the transfer task of Single-wall, we sample the target position within the range of ($\mathcal{U}[3.0,4.0]$, $\mathcal{U}[-1.0,1.0]$).
-Task Single-object Single-wall Multi-wall
-Robot (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$) (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$) (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$)
-Object (1.0, 0.0, 0.0) (1.0, 0.0, 0.0) (1.0, 0.0, 0.0)
-Target (3.0, 0.0) (3.5, 0.0) (4.0, 0.0)
-Wall 1 - (2.0, 0.0, 0.0) (2.0, 0.0, 0.0)
-Wall 2 - - (3.25, 1.25, 0.0)
-Wall 3 - - (3.25, -1.25, 0.0)
-Table B.1: We list the scene configurations, where the initial pose $(x,y,\psi)$ for every existing entity and the target position $(x,y)$ for each task (Single-object, Single-wall, Multi-wall) is shown.
-Appendix C Rewards and Coefficients
-We list the reward functions and corresponding reward parameters for all tasks and training methods in Table C.1.
-The RND reward term is defined as $r_{RND}=\lVert f(s)-\hat{f}(s)\rVert_{2}$, where $s$ is the state to be encoded, and $f(s)$ and $\hat{f}(s)$ are the target and predictor networks, respectively.
+TABLE I: Detailed observation space for each module.
+IV-B High-level Planning Policy
+The high-level policy $\bm{\pi}_{h}$ is trained to generate desired robot velocity commands $\bm{a}_{h}:=\bm{v}_{r,des}$ that achieve successful task completion in object pushing.
+Its observation $\bm{o}_{h}$ includes the robot’s linear and angular velocities, the poses of the robot and object, the target object position $\bm{p}_{t}=(x_{p},y_{p})\in\mathbb{R}^{2}$, and the previous action—all expressed in the world frame.
+In scenarios with obstacles (e.g., walls), $\bm{o}_{h}$ also includes their poses in the world frame.
+The complete high-level observation space $\bm{o}_{h}$ is shown in Table I.
+We use Proximal Policy Optimization (PPO) [schulman2017proximal] as the base RL algorithm, with a simple reward function composed of three terms:
+$r=w_{1}\mathds{1}_{\left\lVert\bm{p}_{o}-\bm{p}_{t}\right\rVert_{2}<\epsilon}+w_{2}r_{\text{CAI}}+w_{3}\left\lVert\bm{a}_{h}-\bm{a}_{h,prev}\right\rVert_{2}^{2},$ (2)
+where $\bm{p}_{o}$, $\bm{p}_{t}$ are the current and target object positions, $\epsilon$ is a success threshold, and $\bm{a}_{h}$, $\bm{a}_{h,prev}$ are the current and previous high-level actions, respectively.
+The exploration bonus $r_{\text{CAI}}$ is derived from the CAI measure $C^{j}$ from (1), where we choose the object to be the entity of interest $j$.
+Resorting to an approximation $\tilde{C}_{j}$, we compute the reward $r_{\text{CAI}}$ at a given state $S=\bm{s}$ as
+$r_{\text{CAI}}=\tilde{C}_{j=\text{object}}(s)=$ (3)
+$\frac{1}{K}\sum_{i=1}^{K}D_{\mathrm{KL}}\left(P_{S_{j}^{\prime}\mid S=s,A=a^{(i)}}\bigg\|\frac{1}{K}\sum_{k=1}^{K}P_{S_{j}^{\prime}\mid S=s,A=a^{(k)}}\right),$
+given $K$ actions $\{a^{(i)}\}_{i=1}^{K}$ sampled from the policy.
+In this work, we find it sufficient to use $K=64$.
+This approximation estimates the marginal $P_{S_{j}^{\prime}\mid s}$ using Monte Carlo sampling.
+We model the transition distribution $P_{S_{j}^{\prime}\mid S=s,A=a}$ as a fixed-variance Gaussian (details in Section IV-C), which enables closed-form KL divergence calculation using the Gaussian mixture approximation [durrieu2012lower].
+The CAI reward encourages the agent to reach states where it exerts greater influence over the object, thereby promoting task-relevant exploration and learning.
+The weight $w_{2}$ for the CAI reward scales with the raw CAI score:
+$w_{2}=w_{2,b}+\max\left(0,(r_{\text{CAI}}-\alpha_{1})/\alpha_{2}\right),$ (4)
+where $\alpha_{1}$ is the threshold for scaling and $\alpha_{2}$ controls the scaling rate.
+Finally, we foster more directed exploration by injecting time-correlated noise into the action sampling process during training [eberhard2023pink, hollenstein2024colored].
+Sampling from time-correlated actions reduces the possibility of meaningless back-and-forth behavior that could result from commonly used white-noise samples.
+We select a correlation strength parameterized as $\beta=0.5$, corresponding to a colored noise between white and pink.
+IV-C Dynamics learning
+To calculate the exploration bonus $r_{\text{CAI}}$ in (3), we learn the transition model $P_{S_{j}^{\prime}\mid S=s,A=a^{(k)}}$ for all entities $j$.
+In our current setting, we focus on a single entity—the pushable object—but the framework naturally extends to multiple entities.
+Concretely, this reduces to learning the object transition model $P_{\bm{s}^{\prime}_{o}\mid\bm{s}_{h},\bm{a}_{h}}$.
+The object state $\bm{s}_{o}$ is defined as the object’s position, while the high-level state $\bm{s}_{h}=(\bm{\xi}_{r},\bm{\xi}_{o},\bm{v}_{r},\bm{v}_{o})$ includes the robot’s pose $\bm{\xi}_{r}$ and velocity $\bm{v}_{r}$ and the object’s pose $\bm{\xi}_{o}$ and velocity $\bm{v}_{o}$.
+The high-level action $\bm{a}_{h}$ corresponds to the desired robot velocity, and the next object state $\bm{s}^{\prime}_{o}=\bm{p}^{\prime}_{o}=(x^{\prime}_{o},y^{\prime}_{o})$ denotes its position at the subsequent timestep.
+Instead of using a pretrained model, CAIMAN leverages data collected from high-level interactions during training to efficiently learn the dynamics.
+As described earlier, we model the object’s transition probability $P_{\bm{s}^{\prime}_{o}\mid\bm{s}_{h},\bm{a}_{h}}$ as a fixed-variance Gaussian distribution $\mathcal{N}(\bm{s^{\prime}}_{o};f_{\theta}(\bm{s}_{h},\bm{a}_{h}),\sigma^{2})$, where $f_{\theta}$ is a neural network that predicts the mean of the distribution.
+To enhance learning efficiency, we incorporate a simple kinematic prior model $f_{\text{kin}}$, which estimates the next object position $\tilde{\bm{p}}^{\prime}_{o}$ using geometric reasoning based on the relative pose between robot and object and the commanded velocity.
+This estimate is computed by projecting the robot’s velocity $\bm{a}_{h}$ onto the direction from the robot to the object and updating the object’s position accordingly:
+[…]
+Here, $\delta\hat{\bm{p}}=(\bm{p}_{o}-\bm{p}_{r})/\left\lVert\bm{p}_{o}-\bm{p}_{r}\right\rVert_{2}$ is the unit vector pointing from the robot to the object, $\delta t$ is the high-level control step size, and $\epsilon_{p}$ is a distance threshold.
+The conditions ensure that the robot is close enough to and moving toward the object.
+We combine the kinematic prior with a learned residual model $f_{\text{res}}$, parameterized by $\theta$, to capture complex physical interactions beyond the kinematics model, including nonlinear effects such as friction, drag, and collisions with obstacles.
+The final dynamics model is thus:
+$f_{\theta}(\bm{s}_{h},\bm{a}_{h})=f_{\text{kin}}(\bm{s}_{h},\bm{a}_{h})+f_{\text{res}}(\bm{s}_{h},\bm{a}_{h};\theta),$ (6)
+where $f_{\text{kin}}$ is deterministic and independent of $\theta$. We train the residual model by minimizing the mean squared error between the predicted and true object positions:
+$\mathcal{L}_{\text{res}}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\lVert\tilde{\bm{s}}^{\prime(i)}_{o}-\bm{s}^{\prime(i)}_{o}\right\rVert_{2}^{2},$ (7)
+where $\tilde{\bm{s}}^{\prime(i)}_{o}=f_{\text{kin}}(\bm{s}^{(i)}_{h},\bm{a}^{(i)}_{h})+f_{\text{res}}(\bm{s}^{(i)}_{h},\bm{a}^{(i)}_{h};\theta)$.
+Although the high-level action $\bm{a}_{h}$ could, in principle, be sampled from the full support of the desired velocity $\bm{v}_{r,des}$ [seitzer2021causal], not all commands are feasible for the low-level controller to execute under the robot’s current state.
+Therefore, we define $\bm{a}_{h}=\tilde{\bm{v}}_{r}$ to be the achievable velocity, and use $\tilde{\bm{v}}_{r}$ for both dynamics learning and CAI computation.
+Given training samples $\mathcal{D}={(\bm{s}_{h}^{(i)},\bm{a}_{h}^{(i)},\bm{s}^{\prime(i)}_{h})}$, we have access to the achieved robot velocity $\bm{v}^{\prime}_{r}$ at the next timestep.
+When computing CAI as an exploration reward, we assume that the robot’s velocity can only change within a limited range over one high-level step.
+Thus, for any state $\bm{s}_{h}$, we sample the action $\bm{a}_{h}$ from a bounded range centered on the current velocity:
+$\bm{a}_{h}=\tilde{\bm{v}}_{r}\sim\mathcal{U}[\bm{v}_{r}\pm\delta\bm{v}_{r}]$ (8)
+where $\delta\bm{v}_{r}$ is a fixed range of velocity deviation. The limits were empirically determined to be $(\delta v_{r}^{x},\delta v_{r}^{y},\delta\omega_{r}^{z})=(0.3,0.3,0.4)$.
+Figure 4: The Single Object (left), Single Wall (middle), and Multi-Wall (right) tasks. The yellow sphere denotes the object’s target position.
+V EXPERIMENTS
+We trained both high- and low-level policies using Isaac Sim [mittal2023orbit].
+The low-level policy operates with a control interval of 0.02 seconds, while the high-level policy has a control interval of 0.2 seconds.
+At each iteration of the high-level control loop, we advance the environment by 0.2 seconds, compute rewards and terminations, collect transitions for both the high-level and dynamics policies, and augment the rewards buffer with the CAI exploration bonus.
+We evaluated CAIMAN on three pushing tasks of increasing difficulty, illustrated in Figure 4.
+In the Single-object task, the robot pushes a single cuboid object to a fixed target position.
+The Single-wall task introduces a fixed wall that blocks the direct path to the target, requiring the robot to maneuver the object around the obstacle.
+The Multi-wall task further increases complexity with multiple fixed walls through which the robot must navigate the object.
+The wall obstacles in the single-wall and multi-wall tasks introduce regions in the state space where the object’s dynamics become more complex, thereby reducing the robot’s ability to exert consistent influence over the object.
+We trained 1500 iterations for the single object and single wall tasks, and 2000 iterations for the multi-wall task, where each iteration consists of 10 high-level control steps.
+All tasks have an episode duration of 20 seconds and use a fixed target position.
 Reward functions
@@
 Multi-wall 40 -5e-3 10 0.01 40 12e-5 1.5e-6
-Table C.1: Simulation training reward parameters.
+TABLE II: Simulation training reward parameters.
 All CAI methods (CAI-kinematic, CAI-learned, CAIMAN) use the CAI reward function, all RND methods (RND-full, RND-object) use the RND reward function, while only the Heuristics method uses the heuristics reward function.
-Appendix D Additional training details
-The low-level policy operates with a control interval of 0.02 seconds, while the high-level policy has a control interval of 0.2 seconds.
-For every iteration through the high-level control loop, we step the environment by 0.2 seconds, compute the rewards and terminations, collect transitions for the high-level and dynamics policies, and calculate and add the CAI exploration bonus to the rewards buffer.
-We trained 1500 iterations for the single object and single wall tasks and 2000 iterations for the multi-wall task, where each iteration consists of 10 high-level control steps.
-All tasks have an episode length of 20 seconds.
-In addition to CAI, we also generate meaningful exploration through colored noise during policy rolling out [68, 69].
-By sampling from time-correlated actions, we reduce the possibility of meaningless back-and-forth behavior that could result from commonly used white-noise samples.
-We select a correlation strength parameterized as $\beta=0.5$, corresponding to a colored noise between white and pink.
-We list all the hyperparameters used for training the high-level loco-manipulation policy and dynamics residual model in Table D.1.
-Environment hyperparameters
-number of environments 4096
-$v_{r,des}^{x}$ (m/s) $[-1.0,1.0]$
-$v_{r,des}^{y}$ (m/s) $[-1.0,1.0]$
-$\omega_{r,des}^{z}$ (rad/s) $[-1.0,1.0]$
-$\delta{v}_{r}^{x}$ (m/s) $0.3$
-$\delta{v}_{r}^{y}$ (m/s) $0.3$
-$\delta{\omega}_{r}^{z}$ (rad/s) $0.4$
-threshold for kinematic prior model $\epsilon_{p}$ (m) $0.7$
-threshold for sparse task reward $\epsilon$ (m) $0.1$
-success rate threshold $\epsilon_{s}$ (m) $0.15$
-high-level policy frequency (Hz) $5$
-low-level policy frequency (Hz) $50$
-number of sampled action in CAI $K$ 64
-PPO hyperparameters
-policy network $[512,256,128]$
-policy activation function ELU
-policy initial standard deviation $1.0$
-value network $[512,256,128]$
-value activation function ELU
-number of mini-batches $4$
-number of epochs per iteration $5$
-clip threshold $0.2$
-learning rate schedule adaptive
-desired KL divergence $0.01$
-value function loss coefficient $1.0$
-entropy coefficient $0.005$
-max gradient norm $1.0$
-$\gamma$ $0.99$
-$\lambda$ $0.95$
-Dynamics hyperparameters
-fixed variance $\sigma$ $1.0$
-residual model $[256,128]$
-batch size $4096$
-number of epochs per iteration $8$
-learning rate $0.0001$
-buffer size $10000000$
-Table D.1: High-level policy and dynamics residual training hyperparameters.
-Appendix E Domain randomization for hardware deployment
-Table E.1 presents the environment parameters that were randomized in simulation during training policies for hardware deployment.
-Parameter Range
-Object mass $(kg)$ $\mathcal{U}[3.5,7.5]$
-Object friction coefficient $\mathcal{U}[0.5,1.5]$
-Initial robot position $(m)$ x: $\mathcal{U}[-0.1,0.1]$, y: $\mathcal{U}[-0.1,0.1]$
-Initial object position $(m)$ x: $\mathcal{U}[0.9,1.1]$, y: $\mathcal{U}[-0.1,0.1]$
-Table E.1: Randomized parameters in simulation for training HL policies for hardware.
-Appendix F Additional Experiment: Dense task reward
-Though we primarily focused on a sparse task reward, we also trained all tasks in simulation under a dense task reward for all methods.
-The dense task reward, defined as $r_{\text{task}}^{dense}=\exp{-\left\lVert\bm{p}_{o}-\bm{p}_{t}\right\rVert_{2}}$, awards an amount inversely proportional to the euclidean distance between the object and the target.
-Combining the dense task rewards with a distance-based heuristic reward resembles the settings from prior work [25].
-Task performance during training is shown in Figure F.1.
-For learning the single object task with dense rewards, all methods demonstrate similar learning speeds and success rates, which indicates that within simpler, less cluttered environments, a dense reward scheme sufficiently guides task learning.
-However, for the dense reward single wall and multi-wall tasks, all CAI-imbued learning methods (CAIMAN, CAI-kinematic, CAI-learned) exhibit better sample efficiency over the heuristics method, indicating that learning the actions necessary for obstacle navigation is facilitated through CAI exploration.
-CAIMAN and CAI-kinematic achieve the fastest convergence and the highest success rates, while CAI-learned is slightly less sample efficient.
-This can be attributed to the fact that learning dynamics from scratch initially produces erroneous predictions, which results in inaccurately high CAI scores and diverts the robot from acquiring task specific skills.
-In contrast, with only the kinematic prior model, the resulting exploration behavior is less adversely impacted when skill learning is heavily guided by dense rewards.
-HeuristicsCAI-kinematicCAI-learnedCAIMAN
-Figure F.1: Policy success rate during training under a dense task reward for all methods. All experiments utilize a minimum of 2 seeds.
-Appendix G Additional Experiment: Cart
-To demonstrate CAIMAN’s effectiveness in handling complex objects with asymmetric dynamics, we provide an additional experiment: pushing a simplified cart with two normal and two caster wheels, as shown in Figure G.1.
-The scene configuration is identical to that of the Single object task, with the box being replaced by the cart.
-Our results demonstrate that CAIMAN is more sample efficient and achieves a higher asymptotic performance (approx.
+We compared our approach against several baselines.
+The Heuristics baseline trains the high-level policy using a distance-based heuristic reward, $r_{\text{heu}}=\exp\left(-\left\lVert\bm{p}_{r}-\bm{p}_{o}\right\rVert_{2}\right)$, which encourages the robot to minimize its distance to the object and mirrors the primary exploration reward used in prior work [jeon2023learning].
+We further evaluated CAIMAN against intrinsic motivation algorithms, specifically Random Network Distillation (RND) [burda2018exploration], implementing two variants: RND-full, which follows the original formulation and computes prediction error over the full state, and RND-object, which restricts prediction error to the object state only.
+For ablation studies, we introduced CAI-kinematic, where CAI is computed using only the kinematic prior (Equation 5), and CAI-learned, where CAI relies on a fully learned dynamics model (no prior).
+The reward functions and corresponding parameters for all tasks and training methods are summarized in Table II.
+The RND reward term is given as $r_{RND}=\lVert f(s)-\hat{f}(s)\rVert_{2}$, where $s$ is the state to be encoded, and $f(s)$ and $\hat{f}(s)$ denote the target and predictor networks, respectively.
+HeuristicsCAI-kinematicCAI-learnedRND-fullRND-objectCAIMAN
+Figure 5: Policy success rate evaluated at every 100 training iterations for all methods and tasks.
+Results are evaluated across 800 episodes and averaged over 3 seeds, shaded area represents standard deviation.
+V-A Simulation results
+We evaluated the learning performance of each method by measuring the success rate for each task.
+A task is successfully completed if the distance between the object and the target is below a threshold $\epsilon_{s}$ at the end of an episode.
+The results are shown in Figure 5.
+With only sparse task rewards, task-relevant skill acquisition heavily relies on effective exploration.
+In the Single-object task, all methods—except RND-full—achieve some degree of success.
+CAIMAN, RND-object, and Heuristics perform the best, while CAI-kinematic and CAI-learned underperform.
+The low performance of CAI-learned is likely due to the high data and time demands required to learn an accurate dynamics model from scratch.
+In contrast, CAI-kinematic performs comparably with the top methods, suggesting that the naive kinematic prior is sufficient for simpler tasks with straightforward object interactions.
+In the more complex Single-wall and Multi-wall tasks, only CAIMAN and RND-object demonstrate notable success, with RND-object showing lower asymptotic performance than CAIMAN in both scenarios.
+This highlights CAIMAN’s unique ability to guide task-relevant exploration even in the presence of obstacles, validating its effectiveness in cluttered environments and independence from dense extrinsic rewards.
+The failure of CAI-kinematic and CAI-learned in these tasks emphasizes the importance of combining a kinematic prior with a learned residual model to efficiently capture the complex object dynamics needed for effective exploration.
+RND-full results in agents randomly exploring the state space without yielding meaningful behavior in all tasks.
+This finding reinforces the need for object-centric exploration bonuses when the skill to be learned involves physical interaction with objects.
+The performance gap between RND-object and CAIMAN can be partially attributed to the phenomenon of detachment [ecoffet2019go], where the agent drifts away from reward-depleted areas and continuously pursues new regions with higher intrinsic reward, potentially losing avenues of task-relevant exploration.
+To validate CAI as an effective exploration signal, Figure 6 explicitly plots the raw CAI reward alongside robot-object interactions in the Single-object task.
+Elevated CAI values during contact states empirically confirm the role of CAI in guiding exploration.
+CAI rewardcontact
+Figure 6: Scaled CAI reward overlaid with a binary indicator of robot–object contact (1-contact, 0-no contact) over one episode of the Single-object task.
+Representative environment frames are shown above.
+V-B Loco-manipulation of irregular objects
+To demonstrate CAIMAN’s effectiveness in handling complex objects with asymmetric dynamics, we conducted an additional experiment using a simplified cart with two normal and two caster wheels, as illustrated in Figure 7.
+The scene configuration is identical to that of the Single-object task, with the cuboid being replaced by the cart.
+Our results show that CAIMAN is more sample efficient and achieves a higher asymptotic performance (approx.
 90% success rate) than most baselines and is on par with the best competitor.
-Notably, this experiment did not require any changes to the framework or physics prior, highlighting the capability of the learned model in estimating complex dynamics and the effectiveness of CAIMAN for learning with irregular objects.
-This environment has no obstacles, but we hypothesize that cluttered settings with walls would further emphasize CAIMAN’s advantages over baselines.
+In particular, this experiment did not require any changes to the original framework or to the kinematics prior, which was derived assuming an object with symmetric motion.
+These results highlight the ability of the learned model to accurately capture complex dynamics and confirm CAIMAN’s capability in learning with irregular objects.
+This environment has no obstacles, but we hypothesize that adding walls would further emphasize CAIMAN’s advantages over other baselines.
 HeuristicsCAI-kinematicCAI-learnedRND-fullRND-objectCAIMAN
-Figure G.1: Left: A simplified cart object. Right: Single cart pushing results.
+Figure 7: Left: A simplified cart object. Right: Single cart pushing results.
+V-C Leveraging pretrained dynamics for new tasks
+Since our framework learns the environment dynamics, we propose that the dynamics residual obtained from a previous training can be reused across different tasks—as long as the underlying dynamics remain consistent.
+We hypothesize that leveraging a pretrained model improves the accuracy of CAI estimates in the early stages of training, thus enhancing the exploration signal and significantly boosting sample efficiency.
+To test this hypothesis, we consider a generalization of the Single-wall task in which the target position is randomized instead of fixed, sampled uniformly within a predefined area as illustrated in Figure 8.
+We compare the original CAIMAN, which learns the dynamics residual from scratch, to one that reuses a pretrained residual model.
+For the latter, we initialize the residual with one from the original Single-wall task and continuously fine-tune it during training with data collected from the new generalized task.
+with pretrainedw/o pretrained
+Figure 8: Left: Single wall task with target positions randomly sampled within an area in front of the wall.
+Right: Learning the single wall random target task using models without pretraining (learned from scratch) and models pretrained from the fixed target task.
+As shown in Figure 8, the reuse of learned dynamics significantly accelerates learning with sparse rewards.
+The pretrained model yields more informative CAI estimates, guiding the robot toward meaningful interaction behaviors early in training.
+This leads to substantial gains in sample efficiency, supporting our hypothesis and demonstrating the benefits of transferring learned dynamics across related tasks.
+V-D Hardware deployment
+We validated our trained policies on real-world quadruped pushing tasks with the Unitree Go2, as shown in Figure 9.
+To bridge the sim-to-real gap, we applied domain randomization to the object’s mass and friction during high-level policy training.
+To obtain the observations for the high-level policies, all entities in the environment were tracked using an external motion capture system.
+Our trained policies were directly deployed to the robot and successfully executed pushing tasks without requiring additional fine-tuning.
+For further details and visualizations of our trained policies in simulation and on hardware, please refer to the supplementary video.
+VI CONCLUSION
+We presented CAIMAN, a general framework for training whole-body pushing skills in legged robots.
+CAIMAN adopts a hierarchical control strategy that decouples locomotion from high-level planning and introduces an intrinsic CAI-based exploration bonus, encouraging control over relevant entities without hand-crafted reward shaping.
+The CAI computation is bootstrapped with a simple yet effective kinematic prior that is refined by a learned residual dynamics model.
+Across diverse quadruped pushing tasks, CAIMAN outperforms competitive baselines in sample efficiency, especially in obstacle-rich scenarios, and moreover generalizes to irregular objects.
+We also showed that the learned dynamics residual accelerates training in new tasks, highlighting the transferability of our approach.
+Finally, hardware experiments demonstrate the seamless deployment of policies trained with CAIMAN on a real quadruped.
+Overall, CAIMAN provides a robust and scalable solution for physically grounded manipulation behaviors in legged robots, paving the way for more autonomous and adaptable robotic systems.
+Figure 9: Snapshots from hardware deployment for the single object and single wall tasks. The mass of the box is 5.5 kilograms with a dimension of $(0.55,0.55,0.5)$ meters.